# Steering LVLMs via Sparse Autoencoder for Hallucination Mitigation

### v1

@@ Appendix D Choice of Scaling Factor
$\text{CHAIR}_{S}$downarrow $\text{CHAIR}_{I}$downarrow Avg.Len
@@
w/ SSL $\gamma=0.2$ 38.52 11.45 102.62
w/ SSL $\gamma=0.4$ 10.16 7.90 64.04
w/ SSL $\gamma=0.6$ 0.76 21.99 349.86
@@
Appendix E Reverse-SSL Approach
@@
Appendix F More Qualitative Results
Figures 1314, and 15 present additional qualitative examples on the LLaVA-Next, LLaVA-1.5, and InstructBLIP models, respectively, to demonstrate the effectiveness of our proposed SSL approach in mitigating hallucinated objects. With the integration of SSL, the generated descriptions by LVLMs exhibit improved fidelity to the visual content while maintaining the richness and informativeness of the language output.
Appendix G Details on the GPT-4o Evaluation
To evaluate the performance of LVLMs on the LLaVA-Bench benchmark, we adopt GPT-4o as the reference evaluator. Following the template provided in Table 8 of Gong et al. (2024), each evaluation instance includes the original image, the base output of the LVLM, and its SSL-enhanced counterpart. The evaluation focuses on both the accuracy and fineness of the generated responses. To mitigate potential biases caused by output order, we randomly swap the positions of the two outputs with a probability of 0.5 before each evaluation. Each sample is evaluated four times to compute an average score. Figures 16,  17, and  18 illustrate representative evaluation examples based on three model architectures: LLaVA-Next, LLaVA-1.5, and InstructBLIP, respectively.
Appendix H License

### v2

@@ Appendix E Choice of Scaling Factor
$\text{CHAIR}_{S}$ $\downarrow$ $\text{CHAIR}_{I}$ $\downarrow$ Avg.Len
@@
w/ SSL $\gamma=0.1$ 44.04 12.64 102.86
w/ SSL $\gamma=0.2$ 41.92 12.10 100.26
w/ SSL $\gamma=0.3$ 38.40 11.67 105.38
w/ SSL $\gamma=0.4$ 37.21 11.45 119.39
w/ SSL $\gamma=0.5$ 8.61 10.55 44.49
w/ SSL $\gamma=0.6$ 3.42 14.02 52.14
@@
Appendix F Reverse-SSL Approach
@@
Appendix G More Qualitative Results
Figures 1415, and 16 present additional qualitative examples on the LLaVA-Next, LLaVA-1.5, and InstructBLIP models, respectively, to demonstrate the effectiveness of our proposed SSL approach in mitigating hallucinated objects. With the integration of SSL, the generated descriptions by LVLMs exhibit improved fidelity to the visual content while maintaining the richness and informativeness of the language output.
Appendix H Details on the GPT-4o Evaluation
To evaluate the performance of LVLMs on the LLaVA-Bench benchmark, we adopt GPT-4o as the reference evaluator. Following the template provided in Table 10 of Gong et al. (2024), each evaluation instance includes the original image, the base output of the LVLM, and its SSL-enhanced counterpart. The evaluation focuses on both the accuracy and fineness of the generated responses. To mitigate potential biases caused by output order, we randomly swap the positions of the two outputs with a probability of 0.5 before each evaluation. Each sample is evaluated four times to compute an average score. Figures 17,  18, and  19 illustrate representative evaluation examples based on three model architectures: LLaVA-Next, LLaVA-1.5, and InstructBLIP, respectively.
Appendix I License